# Large Language Models for Scientific Idea Generation: A Creativity-Centered Survey

### v1

@@ Abstract
Scientific idea generation lies at the heart of scientific discovery and has driven human progress—whether by solving unsolved problems or proposing novel hypotheses to explain unknown phenomena. Unlike standard scientific reasoning or general creative generation, idea generation in science is a multi-objective and open-ended task, where the novelty of a contribution is as essential as its empirical soundness. Large language models (LLMs) have recently emerged as promising generators of scientific ideas, capable of producing coherent and factual outputs with surprising intuition and acceptable reasoning, yet their creative capacity remains inconsistent and poorly understood. This survey provides a structured synthesis of methods for LLM-driven scientific ideation, examining how different approaches balance creativity with scientific soundness. We categorize existing methods into five complementary families: External knowledge augmentation, Prompt-based distributional steering, Inference-time scaling, Multi-agent collaboration, and Parameter-level adaptation. To interpret their contributions, we employ two complementary frameworks: Boden’s taxonomy of Combinatorial, Exploratory and Transformational creativity to characterize the level of ideas each family expected to generate, and Rhodes’ 4Ps framework—Person, Process, Press, and Product—to locate the aspect or source of creativity that each method emphasizes. By aligning methodological advances with creativity frameworks, this survey clarifies the state of the field and outlines key directions toward reliable, systematic, and transformative applications of LLMs in scientific discovery.
@@ 1 Introduction
Scientific discovery has long stood at the frontier of human progress, from uncovering the laws of physics to developing transformative medicines. At the heart of this process lies scientific idea generation—the ability to pose novel, plausible hypotheses that can guide subsequent experimentation and theory building. Unlike routine problem solving, scientific ideation must balance two demanding criteria: novelty, which drives genuine innovation, and valueness, which ensures that the ideas generated are correct, feasible, and valuable (Boden2004).
@@
Recent advances in large language models (LLMs) offer an unprecedented opportunity to augment this process. On the valueness side, research in reasoning and factuality has improved the reliability of model outputs. Techniques such as Chain-of-Thought prompting (wei2022chainofthought), Self-Consistency decoding (wang2023selfconsistency), Tree-of-Thoughts (yao2023tree), and iterative refinement methods like Self-Refine (liu2023selfrefine) enhance structured reasoning. Inference-time scaling approaches, from repeated sampling (brown2024largelanguagemonkeysscaling) to advanced search (kim2023montecarlothought) and frontier RL post-training models (deepseekai2025deepseekr1incentivizingreasoningcapability; yin2025scientificreasoningllmstraining), further demonstrate that more compute can unlock deeper reasoning abilities. Complementary strategies including retrieval-augmented generation (RAG) (lewis2020rag) and agentic tool use integrate external knowledge sources and fact-checking pipelines to ensure factual grounding (schick2023toolformer). Collectively, these advances provide a foundation for promoting valueness and correctness in scientific ideation.
@@
On the creativity side, alignment-oriented methods push LLMs toward greater originality. Extensions of direct preference optimization (DPO) such as Creative Preference Optimization (CRPO) (2025crpo) and Diverse Preference Optimization (DivPO) (2025divpo) explicitly inject signals for novelty and diversity into the optimization process. However, such methods often excel in domains like storytelling or open-ended writing, where factual complexity is limited, but struggle to generalize to science, where creativity must remain tightly coupled with empirical soundness.
@@
The convergence of reasoning and generative creativity highlights why scientific idea generation is uniquely challenging: it is inherently a multi-objective problem where both novelty and value must coexist. This dual demand calls for a deeper understanding of creativity itself. Creativity, a higher-order cognitive function of the human mind that LLMs do not yet possess (liu2025advanceschallengesfoundationagents), remains poorly understood within this emerging field. To ground our analysis, we turn to cognitive science, which offers robust frameworks for studying and categorizing creativity—with the hope that insights from this literature can help reveal which cognitive dimensions of creativity are captured by current LLM-based approaches to scientific idea generation and which remain underexplored. Rhodes’ seminal 4Ps framework (rhodes1961analysis) conceptualizes creativity as the interaction among four dimensions: the person (the individual or system generating ideas), the process (the cognitive or algorithmic mechanisms involved), the press (the surrounding environment and context), and the product (the evaluable outcome). While all four are jointly constitutive of creativity, later work emphasizes that the Product is typically the dependent, measurable outcome, whereas Person, Process, and Press act as the primary sources shaping its emergence (kozbelt2010theories; 4pCreativityModel).
Complementing this perspective, Boden (Boden2004) distinguishes three levels of creative output: combinatorial creativity, where new ideas arise from novel recombinations of known concepts; exploratory creativity, which searches structured spaces to uncover new but conceptually coherent possibilities; and transformational creativity, which reshapes or expands the conceptual space itself, enabling paradigm-shifting discoveries.
Building on these foundations, this survey—unlike existing works that primarily emphasize engineering pipelines, agent architectures, or task coverage (zheng2025autonomy; luo2025llm4srsurveylargelanguage; eger2025transforming; anonymous2025llm; emergentmind2024scientific; ren2025scientificagents)—examines how approaches in LLM-driven scientific discovery manage the dual objectives of scientific valueness and novelty. We propose that these objectives can be systematically understood through the lens of output creativity levels (combinatorial, exploratory, transformational) and creativity sources (person, process, press), as illustrated in Figure 3, which maps current methods to their primary sources of creativity. We hope that insights from cognitive science can help identify which dimensions are covered and which remain underexplored. To make this analysis concrete, as shown in Figure 2, we categorize existing methods into five complementary families:
@@
    Knowledge and retrieval augmentation: A central line of work enhances LLMs with curated, diverse, and relational domain knowledge, compensating for the limitations of static pre-training. Adding related literature or factual resources into the model’s context not only grounds outputs in established knowledge and reduces hallucinations, but also acts as a source of creativity infusion from the environment. In Rhodes’ 4Ps, this reflects the press, balancing correctness with novelty derived from recombining known elements. Such methods are most likely to foster combinatorial creativity. We expand on these approaches in Section 2.
@@
    Prompt-based distributional steering: Another family of approaches steers LLMs toward more original ideas through input manipulations. Manipulations like role priming, constraint-based prompting, or structured scaffolds encourage less typical outputs without parameter changes. Rooted mostly in the press, they often yield combinatorial creativity but can also broaden into exploratory forms when prompts guide exploration. Details and more examples appear in Section 3.
@@
    Search and sampling expansions: Inference-time scaling—via iterative refinement or branching exploration — expand the capabilities of LLMs beyond single-pass decoding by enabling dynamic exploration and without additional model training. By systematically enlarging the idea search space and moving beyond obvious hypotheses, these methods increase the chance of exploratory creativity, though correctness depends on strong evaluation signals. This family maps most closely to the process dimension. We discuss them further in Section 4.
@@
Finally, we outline open directions in Section 8. Current methods tend to remain at the combinatorial or exploratory level, with transformational creativity still elusive. Much progress has been made from the “process” and “press” perspectives, yet the “person” and “product” dimensions remain underexplored. For example, shifting from idea-level search to agent-level search, as it is done in recent open-ended methods (zhang2025darwin), may unlock richer person-level creativity. At a deeper level, the auto-regressive training paradigm or llm attention-based architecture itself may impose structural limits on the creative potential of LLMs in science (nagarajan2025rolldicelook). On the product side, the lack of standardized metrics and benchmarks for evaluating generated scientific ideas leaves creativity assessment vague and subjective. Addressing these gaps could move us closer to realizing LLMs as true partners in scientific discovery.
@@ 2 Knowledge Augmentation: Grounding LLMs in External Evidence
Scientific ideation rarely begins from a blank slate. Researchers typically draw upon prior literature, raw data, or domain-specific facts to frame questions and generate hypotheses. LLMs, however, operate with a static knowledge base limited to their pretraining cutoff. To better mimic the way scientists ground their thinking, LLMs need access to up-to-date and domain-relevant information during inference. Adding related literature or factual resources into the model’s context not only grounds outputs in established knowledge, reducing hallucinations, but also acts as a source of creativity infusion from the environment. In the language of creativity frameworks, this corresponds to the press dimension of Rhodes’ 4Ps (rhodes1961analysis), balancing correctness with novelty induced from external knowledge. In this section, we survey two major approaches: Semantic Retrieval, which dynamically incorporates evidence from external corpora, and Relational Retrieval, which provide structured, relational grounding. We then discuss their limitations and open challenges. Figure 1(a) illustrates an example of knowledge augmentation with these two approaches.
@@ 2.1 Semantic Retrieval
Originally designed for open-domain question answering, RAG (lewis2020rag) has become foundational in scientific idea generation. Early systems like PaperQA (lala2023paperqa) and LitLLM (agarwal2024litllm) focused on raw literature access, while newer frameworks like Ideasynth (pu2025ideasynth), Scideator (radensky2024scideator), Nova (hu2024nova), ResearchAgent (baek2024researchagent)) emphasize multi-stage decomposition, critique-and-refine loops, and citation-aware retrieval. Furthermore, systems such as SciPIP (wang2024scipip), SciMON (wang2024scimon), SCI-IDEA (keya2025sciidea), and PaperHelper (yin2025paperhelper) push beyond simple matching by integrating novelty detection, visual scaffolding, and hybrid fine-tuning.
@@ 2.2 Relational Retrieval
KG-based approaches allow LLMs to traverse paths across domains, promoting more creative and meta-level reasoning than local similarity alone. In biomedicine, systems like PubTator (wei2013pubtator; wei2024pubtator30aipoweredliterature) and SciSpacy (neumann2019scispacy) extract subgraphs from literature to enable link prediction and hypothesis generation.
Scenario 1: Similarity-driven vs. Relation-driven retrieval Seed research goal: Design methods to make LLMs more reliable when solving reasoning tasks in out-of-distribution (OOD) settings. Semantic-driven Retrieval. The system generates a query such as “robustness in LLM reasoning under OOD” and retrieves the top-$k$ semantically similar papers from a vector database or literature API. Illustrative retrieved items: Paper A: Chain-of-thought prompting improves reasoning in arithmetic tasks. Paper B: Self-consistency decoding reduces reasoning errors in LLMs. Paper C: Fine-tuning LLMs with synthetic reasoning chains. Generated hypothesis: Combine chain-of-thought prompting with self-consistency decoding and fine-tune on synthetic reasoning datasets to improve robustness on novel tasks. Relation-driven Retrieval. The seed is mapped to nodes such as “LLM reasoning” and “distribution shift” within a knowledge graph of ML research concepts and papers. The system performs multi-hop traversal to find relational paths across subfields. Illustrative graph paths: Distribution shift Causal representation learning Counterfactual data augmentation LLM reasoning errors Formal methods verification Program synthesis validation Robustness Neurosymbolic reasoning Graph neural networks $\scriptstyle{addressed\_by}$$\scriptstyle{combined\_with}$$\scriptstyle{mitigated\_by}$$\scriptstyle{applied\_in}$$\scriptstyle{linked\_to}$$\scriptstyle{integrated\_with}$ Generated hypothesis: Develop a hybrid framework where causal representation learning identifies shift-invariant features, augmented by counterfactual data generation, and pair it with neurosymbolic reasoning modules to improve LLM reasoning robustness under OOD conditions. Similarity-driven retrieval emphasizes local relevance by surfacing closely related techniques, while relational retrieval encourages cross-paradigm creativity, uncovering connections unlikely to emerge from similarity alone. Despite these differences, standard greedy decoding in LLMs generally limits outputs from either approach to the combinatorial level.
An early example of Relational Retrieval is Chain of Ideas (li2024chain) that by explicitly tracking breakthroughs and their dependencies, enabled LLMs to reason over trends rather than isolated documents. This inspired subsequent work that generalized the idea to broader graph formalisms. The KG-CoI framework (xiong2024kgcoi) and Grounding LLM Reasoning with Knowledge Graphs (amayuelas2025grounding) demonstrate that graph-structured multi-hop reasoning improves plausibility and interpretability. Large-scale efforts such as Graph of AI Ideas (GoAI) (gao2025goai) and SciMuse (gu2025scimuse) show how domain-spanning graphs can surface trends, gaps, and opportunities invisible to similarity-based retrieval.
@@ 2.3 Takeaways and Limitations
RAG-style and graph-based retrieval answer complementary needs in LLM-assisted scientific ideation. As shown in Scenario 2.2111To aid intuition, we also include scenario boxes throughout this survey. These illustrative examples are synthetic—not actual system outputs—but are designed to clarify how different methods might operate in practice. By grounding abstract mechanisms in concrete cases, the boxes highlight contrasts between approaches and help readers better visualize potential reasoning and generative steps, without implying empirical results., Similarity-driven RAG gives fast, relevant context — it pulls the nearest prior work and facts so the model can be accurate and up-to-date. Relational retrieval, by contrast, exposes structure and cross-domain paths that are invisible to nearest-neighbour search, enabling multi-hop, meta-level, and multidisciplinary leaps.
Despite these advances, heavy reliance on retrieved content can bias ideas toward well-represented clusters, reinforcing conservative or incremental thinking rather than encouraging novelty. Empirical studies confirm these tendencies: higher retrieval quality often reduces output diversity as models overfit to retrieved passages (lee2024inferencescaling); decoding strategies overweight context at the expense of parametric reasoning (wang2025corag); and when external evidence conflicts with internal knowledge, LLMs exhibit confirmation bias rather than creative divergence (lee2025yourai). Since outputs are constrained by the scope and diversity of the augmented knowledge and models often default to simply stitching together retrieved fragments, knowledge augmentation largely supports combinatorial creativity and it struggles to foster exploratory or transformational creativity.
Addressing these limitations will require hybrid strategies that balance grounding with exploration, incorporate richer modalities, and explicitly encourage divergence—directions we revisit in subsequent sections on prompt-based creativity and inference-time scaling.
Scientific Idea Generation Knowledge Augmentation Similarity-Based Retrieval (wang2024scipip) (wang2024scimon) (yin2025paperhelper) Relational-Based Retrieval (gao2025goai) (amayuelas2025grounding) (gu2025scimuse) Prompt-Driven Methods Persona Priming (zhao2025assessingllms) (liu-etal-2025-persona) (kim2025exploringpersonadependentllmalignment) (wan2025usinggenerativeaipersonas) Constraint-Based Queries (denial_prompting_2024) Structured Prompting (wei2022chainofthought) (wan2023cosp) (o2025sparks) (li2024chain) Multilingual Prompting (wang2025mlprompt) (vatsal2025survey) Multi-Agent Systems Pipeline-Oriented Workflows (ghareeb2025robinmultiagentautomatingscientific) (schmidgall2025agentlaboratoryusingllm) (gottweis2025aicoscientist) Debate-Based Loops (johnson2024virsci) (liang2023MAD) Parameter Adaptation Reinforcement Learning (surina2025algorithmdiscoveryllmsevolutionary) (yin2025scientificreasoningllmstraining) (pan2025medvlmr1incentivizingmedicalreasoning) (liu2025drugimprovergptlargelanguagemodel) (ahmed2025improvingtargetedmoleculegeneration) Supervised Finetuning (xie2023darwinseriesdomainspecific) (prabhakar2025omnisciencedomainspecializedllmscientific) Structured Optimization (chennakesavalu2025aligningtransformerscontinuousfeedback) (gkoumas2024almolalignedlanguagemoleculetranslation)
Figure 2: A four-level taxonomy for scientific idea generation methods using LLMs.

### v2

@@ Abstract
Scientific idea generation is central to discovery, requiring the joint satisfaction of novelty and scientific soundness. Unlike standard reasoning or general creative generation, scientific ideation is inherently open-ended and multi-objective, making its automation particularly challenging. Recent advances in large language models (LLMs) have enabled the generation of coherent and plausible scientific ideas, yet the nature and limits of their creative capabilities remain poorly understood. This survey provides a structured synthesis of methods for LLM-driven scientific ideation, focusing on how different approaches trade off novelty and scientific validity. We organize existing methods into five complementary families: External knowledge augmentation, Prompt-based distributional steering, Inference-time scaling, Multi-agent collaboration, and Parameter-level adaptation. To interpret their contributions, we adopt two complementary creativity frameworks: Boden’s taxonomy to characterize the expected level of creative novelty, and Rhodes’ 4Ps framework to analyze the aspects or sources of creativity emphasized by each method. By aligning methodological developments with cognitive creativity frameworks, this survey clarifies the evaluation landscape and identifies key challenges and directions for reliable and systematic LLM-based scientific discovery.
@@ 1 Introduction
Scientific discovery has long stood at the frontier of human progress, from uncovering the laws of physics to developing transformative medicines. At the heart of this process lies scientific idea generation—the ability to propose hypotheses that are not only new but also scientifically meaningful. In cognitive science, this ability is captured by creativity, which defines as requiring two jointly necessary components: novelty and valueness(Boden, 2004). Novelty refers to the degree to which an idea departs from existing knowledge or established patterns, while valueness concerns whether the idea is correct, plausible, feasible, or useful within a scientific context. Novelty without value results in arbitrary or incoherent ideas; value without novelty yields incremental or redundant contributions. Scientific ideation, therefore, emerges only when these two criteria coexist.
@@
Recent advances in large language models (LLMs) offer an unprecedented opportunity to augment this process. On the valueness side, research in reasoning and factuality has improved model reliability. Techniques such as Chain-of-Thought prompting (Wei et al., 2022), Self-Consistency decoding (Wang et al., 2023c), Tree-of-Thoughts (Yao et al., 2023), and iterative refinement methods like Self-Refine (Madaan et al., 2023) enhance structured reasoning. Inference-time scaling approaches, from repeated sampling (Brown et al., 2024) to advanced search (Sprueill et al., 2023) and frontier RL post-training models (DeepSeek-AI et al., 2025; Yin et al., 2025b), further demonstrate that compute can unlock deeper reasoning abilities. Complementary strategies including retrieval-augmented generation (RAG) (Lewis et al., 2020) and agentic tool use integrate external knowledge sources and fact-checking pipelines to ensure factual grounding (Schick et al., 2023). Collectively, these advances provide a foundation for promoting valueness and correctness in scientific ideation.
@@
In contrast, efforts explicitly targeting novelty remain more limited and domain-dependent. Alignment-oriented methods push LLMs toward greater originality. Extensions of direct preference optimization (DPO) such as Creative Preference Optimization (CRPO) (Ismayilzada et al., 2025a) and Diverse Preference Optimization (DivPO) (Lanchantin et al., 2025) explicitly inject signals for novelty and diversity into the optimization process. However, such methods often excel in domains like storytelling or open-ended writing, where factual complexity is limited, but struggle to generalize to science, where creativity must remain tightly coupled with empirical soundness.
The convergence of reasoning and generative creativity highlights why scientific idea generation is uniquely challenging: it is inherently a multi-objective problem where calls for a deeper understanding of creativity itself. Creativity, a higher-order cognitive function of the human mind that LLMs do not yet possess (Liu et al., 2025a), remains poorly understood within this emerging field.
@@
To ground our analysis, we turn to cognitive science, which offers robust frameworks for studying and categorizing creativity and helps clarify which cognitive dimensions are captured by current LLM-based approaches and which remain underexplored. Rhodes’ seminal 4Ps framework (Rhodes, 1961) conceptualizes creativity as the interaction among four dimensions: the person (the individual or system generating ideas), the process (the cognitive or algorithmic mechanisms involved), the press (the surrounding environment and context), and the product (the evaluable outcome). While all four are jointly constitutive of creativity, later work emphasizes that the product is typically the dependent, measurable outcome, whereas person, process, and press act as the primary sources shaping its emergence (Kozbelt et al., 2010; Gruszka and Tang, 2017).
Complementing this perspective, Boden (Boden, 2004) distinguishes three levels of creative output: combinatorial creativity, arising from novel recombinations of existing concepts; exploratory creativity, which results from structured search within a given conceptual space; and transformational creativity, which alters or expands the space itself. In this survey, we use Boden’s taxonomy to characterize the expected creative potential of LLM-based scientific ideation methods according to their dominant mechanisms. Methods centered on recombination are thus expected to yield combinatorial creativity, those enabling systematic exploration to favor exploratory creativity, and those that challenge established assumptions to potentially approach transformational creativity. These assignments are conceptual ,reflecting tendencies rather than guaranteed outcomes.
Building on these foundations, this survey—unlike existing works that primarily emphasize engineering pipelines, agent architectures, or task coverage (Zheng et al., 2025; Luo et al., 2025; Eger et al., 2025; Han and Cheng, 2025; EmergentMind Research Group, 2024; Ren et al., 2025)—examines how approaches in LLM-driven scientific discovery manage the dual objectives of scientific valueness and novelty. We propose that these objectives can be systematically understood through the lens of output creativity levels (combinatorial, exploratory, transformational) and creativity sources (person, process, press), as illustrated in Figure 11, which maps current methods to their primary sources of creativity. Although Boden’s and Rhodes’ frameworks are classical, their intentionally high-level and orthogonal dimensions allow them to map naturally onto modern LLM-based systems. The 4Ps emerged empirically from analyses of creativity definitions and were independently rediscovered across multiple studies, supporting their robustness as an organizing framework (Jordanous, 2012). Unlike alternative creativity-level frameworks such as the 4C model (Kaufman and Beghetto, 2009), whose categories span everyday and personal creativity (e.g., mini-c and little-c), Boden’s taxonomy is tailored to research-level idea generation, where novelty arises through recombination, structured exploration, or transformation of conceptual spaces; as a result, these frameworks continue to serve as foundational baselines in recent studies of LLM creativity (Nagarajan et al., 2025; Gu et al., 2025; Jackson et al., 2025; Ismayilzada et al., 2025b).
To make this analysis concrete, as illustrated in Figure 3, we categorize existing methods into five complementary families according to their primary mechanisms for influencing novelty and valueness in LLM-based ideation:
@@
    Knowledge and retrieval augmentation: A central line of work enhances LLMs with relational domain knowledge, compensating for the limitations of static pre-training. Adding related literature or factual resources into the model’s context not only grounds outputs in established knowledge and reduces hallucinations, but also acts as a source of creativity infusion from the environment. In 4Ps, this reflects the press, balancing correctness with novelty derived from recombining known elements. Such methods are most likely to foster combinatorial creativity. We expand on these approaches in Section 2.
@@
    Prompt-based distributional steering: Another family of approaches steers LLMs toward more original ideas through input manipulations. Manipulations like role priming, constraint-based prompting, or structured scaffolds encourage less typical outputs without parameter changes. Rooted mostly in the press, they often yield combinatorial creativity but can also broaden into exploratory forms when prompts guide a systematic exploration. Details and more examples appear in Section 3.
@@
    Search and sampling expansions: Inference-time scaling—via iterative refinement or branching exploration — expand the capabilities of LLMs beyond single-pass decoding by enabling dynamic exploration and without additional model training. By systematically enlarging the idea search space, these methods increase the chance of exploratory creativity, though correctness depends on strong evaluation signals. This family maps most closely to the process dimension. We discuss them further in Section 4.
@@
This taxonomy is derived by synthesizing recurring design patterns in the LLM-based scientific discovery literature and is intended to be general and extensible rather than exhaustive; hybrid and future methods may span or extend multiple categories. Table 3 summarizes how the five method categories are empirically observed or theoretically expected to influence the creativity.
@@
Finally, we outline open directions in Section 9. Current methods tend to remain at the combinatorial or exploratory level, with transformational creativity still elusive. Much progress has been made from the “process” and “press” perspectives, yet the “person” and “product” dimensions remain underexplored. For example, shifting from idea-level search to agent-level search, as it is done in recent open-ended methods (Zhang et al., 2025a), may unlock richer person-level creativity. At a deeper level, the auto-regressive training paradigm or llm attention-based architecture itself may impose structural limits on the creative potential of LLMs in science (Nagarajan et al., 2025). On the product side, the lack of standardized metrics and benchmarks for evaluating generated scientific ideas leaves creativity assessment vague and subjective. Addressing these gaps could move us closer to realizing LLMs as true partners in scientific discovery.
@@ 2 Knowledge Augmentation: Grounding LLMs in External Evidence
Scientific ideation rarely starts from a blank slate; researchers typically draw on prior literature, data, and domain knowledge to frame questions and generate hypotheses. In contrast, LLMs rely on a static knowledge base limited to their pretraining cutoff, motivating the need for access to up-to-date and domain-relevant information at inference time. Incorporating external literature or factual resources into the model’s context grounds generation in established knowledge, reduces hallucinations, and serves as a source of creative input from the environment. From the perspective of Boden’s taxonomy, knowledge augmentation primarily supports combinatorial creativity by enabling novel ideas through the recombination of existing concepts rather than transformation of the conceptual space. Indeed, Gu et al. demonstrate that retrieval-augmented models can explicitly enable combinatorial creativity by synthesizing research ideas across different abstraction levels. This mechanism aligns with the press dimension of Rhodes’ 4Ps (Rhodes, 1961), where external context introduces novelty while preserving grounding and correctness. We survey two main approaches—Semantic retrieval, which dynamically incorporates evidence from external corpora, and Relational retrieval, which provides structured grounding—and discuss their limitations and open challenges. Figure 1(a) illustrates an example of knowledge augmentation with these two approaches.
@@ 2.1 Semantic Retrieval
Originally designed for open-domain question answering, RAG (Lewis et al., 2020) has become foundational in scientific idea generation. Early systems like PaperQA (Lála et al., 2023) and LitLLM (Agarwal et al., 2024) focused on raw literature access, while newer frameworks like Ideasynth (Pu et al., 2025), Scideator (Radensky et al., 2024), Nova (Hu et al., 2024), ResearchAgent (Baek et al., 2024)) emphasize multi-stage decomposition, critique-and-refine loops, and citation-aware retrieval. Furthermore, systems such as SciPIP (Wang et al., 2024c), SCI-IDEA (Keya et al., 2025), and PaperHelper (Yin et al., 2025a) push beyond simple matching by integrating novelty detection, visual scaffolding, and hybrid fine-tuning.
@@ 2.2 Relational Retrieval
KG-based approaches allow LLMs to traverse paths across domains, promoting more creative and meta-level reasoning than local similarity alone. In biomedicine, systems like PubTator (Wei et al., 2013; 2024) and SciSpacy (Neumann et al., 2019) extract subgraphs from literature to enable link prediction and hypothesis generation.
Figure 2: Comparison of similarity-driven and relation-driven retrieval methods.
An early example of Relational Retrieval is Chain of Ideas (Li et al., 2024a) that by explicitly tracking breakthroughs and their dependencies, enabled LLMs to reason over trends rather than isolated documents. This inspired subsequent work that generalized the idea to broader graph formalisms. The KG-CoI framework (Xiong et al., 2024),SciMON (Wang et al., 2024a) and Grounding LLM Reasoning with Knowledge Graphs (Amayuelas et al., 2025) demonstrate that graph-structured multi-hop reasoning improves plausibility and interpretability. Large-scale efforts such as Graph of AI Ideas (GoAI) (Gao et al., 2025b) and SciMuse (Gu and Krenn, 2025) show how domain-spanning graphs can surface trends, gaps, and opportunities invisible to similarity-based retrieval.
@@ 2.3 Takeaways and Limitations
RAG-style and graph-based retrieval answer complementary needs in LLM-assisted scientific ideation. As shown in Scenario 2111To aid intuition, we also include scenario boxes throughout this survey. These illustrative examples are synthetic—not actual system outputs—but are designed to clarify how different methods might operate in practice. By grounding abstract mechanisms in concrete cases, the boxes highlight contrasts between approaches and help readers better visualize potential reasoning and generative steps, without implying empirical results., Similarity-driven RAG gives fast, relevant context — it pulls the nearest prior work and facts so the model can be accurate and up-to-date. Relational retrieval, by contrast, exposes structure and cross-domain paths that are invisible to nearest-neighbour search, enabling multi-hop, meta-level, and multidisciplinary leaps.
Empirical evidence from GoAI study— which explicitly organizes and traverses ideas using a structured graph— (Gao et al., 2025b) shows that GoAI achieves higher Elo-based novelty scores (3.83) and the highest overall average score (3.39) compared to less structured baselines, such as Chain-of-Ideas(Li et al., 2024a) (3.21 novelty, 3.27 average) and vanilla LLM prompting (2.44 novelty, 2.70 average). These results indicate that relational structuring of ideas does more than improve relevance or clarity: it supports combinatorial creativity by enabling non-trivial recombination across ideas, while maintaining feasibility and effectiveness 3.
Despite these advances, heavy reliance on retrieved content can bias ideas toward well-represented clusters, reinforcing conservative or incremental thinking rather than encouraging novelty. Empirical studies confirm these tendencies: higher retrieval quality often reduces output diversity as models overfit to retrieved passages (Lee et al., 2024); decoding strategies overweight context at the expense of parametric reasoning (Wang et al., 2025b); and when external evidence conflicts with internal knowledge, LLMs exhibit confirmation bias rather than creative divergence (Lee et al., 2025).
Addressing these limitations may require hybrid strategies that balance grounding with exploration, incorporate richer modalities, and explicitly encourage divergence—directions we revisit in subsequent sections on prompt-based creativity and inference-time scaling.
{forest}
Figure 3: A five-level taxonomy for scientific idea generation methods using LLMs.